### v1

@@ Abstract
keywords:
Lattice dynamics, anharmonicity, force constants, elastic properties, phonons
@@
The state of thermal equilibrium (atomic positions and crystal shape and size) can be predicted at any temperature and volume (or pressure)
@@
Can also perform molecular dynamics simulations and calculate atomic trajectories and heat current using the extracted anharmonic force field. This is the more accurate alternative to solving BE, as it treates anharmonicity and mass disorder beyond perturbation theory.
@@ 1 Introduction
The theory of lattice dynamics, based essentially on the physics of the harmonic oscillator has shaped the science behind the thermophysical properties of materials. Nowadays, with the advances in computer software and hardware, calculations of the electronic ground state properties of materials, based on the density functional theory has become routine. Total energy and forces on atoms are an output of such theories, and this has allowed us Esfarjani and Stokes (2008) to develop methodologies to compute force derivatives, or force constants (FCs), in order to obtain the parameters of an anharmonic lattice dynamics model. Many softwares are nowadays available to compute force constants of crystalline solids. Among them, we can name ALAMODE Tadano et al. (2014); Tadano and Tsuneyuki (2018), ShengBTE Li et al. (2014), AlmaBTE Carrete et al. (2017), Phonopy Togo and Tanaka (2015), SSCHA Monacelli et al. (2021) , Hiphive Eriksson et al. (2019), TDEP Hellman et al. (2013); Hellman and Abrikosov (2013), Phonon Parlinski et al. (1997), Phon Alfè (2009) … and many others. Here, we are presenting a new set of codes based on our original paper Esfarjani and Stokes (2008) to compute thermophysical properties of anharmonic crystals. These codes are similar in capabilities but different in the way the FCs are extracted and processed. Our code to calculate the thermal conductivity (THERMACOND) uses full symmetry of the crystal to compute the non-equilibrium distribution function more efficiently. Codes to compute the free energy and the state of equilibrium at high temperatures, as well as the computation of mechanical properties are relatively novel. In addition, we provide codes (ANFOMOD) to perform molecular dynamics simulations and calculate the heat current with the extracted polynomial force field. In what follows, we describe the methodology and capabilities of the code to extract force constants, FOCEX, which stands for FOrce Constants EXtraction. The FCs are extracted from a set of displacement-force data on atoms in one or many supercells. They can be used as input to the code ANFOMOD (ANharmonic FOrce MOlecular Dynamics) to perform MD simulations and calculate the heat current and thermal conductivity from the exact Green-Kubo formula, thereby incorporating anharmonicity and defect scattering beyond perturbation theory. Our other codes, which take as input the extracted force constants, compute the thermal conductivity (THERMACOND) and the state of thermal equilibrium, including structural phase changes, at finite temperatures (SCOP8). They will be described in other upcoming papers.
@@ 2.2 Invariance under arbitrary translations and rotations
The potential energy of a solid remains the same under arbitrary translations and rotations of that solid. For any given atomic displacement $u$, the potential energy associated with the translation of the whole system by an arbitrary vector $c$ is $V(u+c)$ and it must satisfy $V(u+c)=V(u)$ and $F(u+c)=F(u)$, where $u$ are the dynamical variables and $c$ is the arbitrary constant shift vector. Identifying terms order by order we arrive at the following equations Liebfried and Ludwig (1961):
@@
Similarly, the force on atoms and the total potential energy of the crystal should not change under arbitrary rotation of the crystal. This leads to the following constraints on FCs Liebfried and Ludwig (1961):
@@
We see that rotational invariance relations relate the first to the second-order terms, the second to the third-order terms and the third to the fourth-order terms, respectively. This implies that if the expansion of the potential energy is truncated after the fourth order terms, we must start with the fourth order terms, and the application of the rotational invariance rules will give us constraints on the third and lower-order FCs. For more details on the proof regarding the translation and rotational invariance relations, we refer the reader to the article by Liebfried and Ludwig Liebfried and Ludwig (1961).
@@
In addition to these symmetries, it is important to note that elastic constants are subject to a number of constraints that enforces them to be symmetric under the exchange of Voigt indices. Huang discusses these invariancesBorn and Huang (1954), which provide an additional 15 set of constraint equations for low symmetry crystals with internal degrees of freedom in the primitive cell Sluiter et al. (1998, 1999)]. These relations are explicitly written in section 3 where we define the elastic constants.
@@ 2.3 Invariance under point or space group symmetry
Other symmetry operations such as rotations or mirror symmetries, elements of the space group also leave the potential energy of the crystal unaltered Maradudin and Fein (1962). In particular, force constants must be invariant under translations of lattice vectors, implying the following relations:
@@ 2.5 Phonon dispersion and their density of states
$D_{\tau,\tau^{\prime}}^{\alpha\beta}(k)\,e_{\tau^{\prime}\beta,\lambda}(k)=\omega_{\lambda}^{2}(k)\,e_{\tau\alpha,\lambda}(k)\,\,\,\,and\,\,\,\,\sum_{\lambda}e_{\tau\alpha,\lambda}(k)^{*}e_{\tau^{\prime}\beta,\lambda}(k)=\delta_{\alpha\beta}\delta_{\tau\tau^{\prime}}\,\,\,\,and\,\,\,\,\sum_{\tau\alpha}e_{\tau\alpha,\lambda}(k)^{*}e_{\tau\alpha,\lambda^{\prime}}(k)=\delta_{\lambda\lambda^{\prime}}$ (9)
@@
The phonon density of states is calculated using both the tetrahedron method in the full first Brillouin zone (FBZ) and the Gaussian-smearing method in the irreducible Brillouin zone (IBZ). The latter will be inaccurate unless a large number of kpoints are used in the sampling of the IBZ, while the former is almost converged for a moderate grid size on the order of 20 k-mesh per direction.
@@ 2.6 Non-analytic correction for ionic dielectrics
For such systems, the input to the file dielectric.params which contains the electronic contribution to the dielectric matrix $\epsilon_{\infty}$ and the Born effective charges $Z_{\tau}^{\alpha\beta}$ of every atom $\tau$ in the primitive cell, must be properly initialized. These numbers are obtained from a separate DFPT calculationGonze and Lee (1997); Baroni et al. (2001). The contribution of the ionic charges on the force constants is long-ranged and cannot be captured by the small supercell. It is usually calculated using the Ewald sum method Ewald (1921); Ohno et al. (2018); Gonze and Lee (1997), which is formed from the the sum of short-range real space component and a long-range reciprocal-space component. In FOCEX, once the harmonic force constants are fitted, we calculate the dynamical matrix on a grid of reciprocal lattice vectors of the supercell inside the Wigner-Seitz cell of the primitive lattice, including its boundaries, albeit with a weight inversely proportional to the number of images that also fall on the Brillouin zone boundaries. This corresponds to the Fourier transform of the real-space force constants. Phonon frequencies on this grid, $D(G_{\rm sc})$, are exact. From these dynamical matrices, we subtract the non-analytical contributionGonze and Lee (1997), and Fourier transform back to real space, in-order to presumably obtain a short-ranged set of force constants, with a range confined to the Wigner-Seitz cell of the supercell. Then at an arbitrary wavevector $q$, we use Fourier interpolation to calculate the dynamical matrix from the latter set of force constants, to which the non-analytical part that was subtracted earlier, is added back (see also Togo et al. (2023)).
@@ 3 Some notes on Elastic constants
It is possible to extract the elastic constants from the knowledge of the force constantsWallace (1998). Below, we outline a simpler method and provide the final formulas.
@@
While this is valid strictly at zero temperature, this discussion can be extended to finite temperatures using the self-consistent phonon theoryWerthamer (1970) where the harmonic force constant $\Phi$ is replaced by an effective temperature-dependent oneEsfarjani and Liang (2020, 2022): $\Phi\rightarrow K(T)=\Phi+\Psi S+\chi(SS+\langle yy\rangle)/2$. In what follows, we will derive the (harmonic) elastic properties, strictly speaking, at zero temperature, but the above formula may be used as an extension to finite temperatures, by replacing every occurrence of $\Phi$ in the following by $K(T)$.
@@
Compared to the method discussed by Wallace Wallace (1998), here, we are proposing an equivalent but simpler method to obtain elastic constants. We first write the harmonic potential energy in eq. 12, phenomenologically, up to second powers of internal coordinates $u^{0}$ and strain $\eta$ as:
$V^{LD}(\eta,u^{0})=E_{0}+\Pi_{\tau\alpha}u^{0}_{\tau\alpha}+\Xi_{\mu\nu}\eta_{\mu\nu}+\frac{1}{2}\left[\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}\,u^{0}_{\tau\alpha}u^{0}_{\tau^{\prime}\alpha^{\prime}}+A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}\,\eta_{\alpha\beta}\eta_{\alpha^{\prime}\beta^{\prime}}+2Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}\,u^{0}_{\tau\alpha}\eta_{\alpha^{\prime}\beta^{\prime}}\right]$ (13)
@@
$\Xi_{\alpha\beta}$ $=$ $\frac{1}{N}\sum_{R\tau}\Pi_{\tau\alpha}(R+\tau)^{\beta}=\sum_{\tau}\Pi_{\tau\alpha}\tau^{\beta}$ (14)
$\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}$ $=$ $\frac{1}{N}\sum_{R,R^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}\,\,\,\,{\rm with}\,\,\,\,\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\phi_{\tau^{\prime},\tau}^{\alpha^{\prime}\alpha};\,\,\,\,\,\,{\rm ASR}=>\sum_{\tau}\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\sum_{\tau^{\prime}}\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=0$ (15)
$Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}$ $=$ $\frac{1}{N}\sum_{RR^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R^{\prime}+\tau^{\prime})^{\beta^{\prime}}=\frac{1}{N}\sum_{RR^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R^{\prime}+\tau^{\prime}-R-\tau)^{\beta^{\prime}}\,\,\,\,{\rm ASR}=>\sum_{\tau}Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}=0$ (16)
$A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}$ $=$ $\frac{1}{N}\sum_{R\tau,R^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R+\tau)^{\beta}(R^{\prime}+\tau^{\prime})^{\beta^{\prime}}\,\,\,\,{\rm with}\,\,\,\,A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}=A_{\alpha^{\prime}\beta^{\prime},\alpha\beta}$ (17)
@@
$u^{0}_{\tau\alpha}(\eta)=-\Gamma_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}(\Pi_{\tau^{\prime}\alpha^{\prime}}+Q_{\tau^{\prime}\alpha^{\prime};\mu\nu}\,\eta_{\mu\nu});\,\,{\rm with}\,\,\,\Gamma_{1,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\Gamma_{\tau,1}^{\alpha\alpha^{\prime}}=0\,\,\,\,\,\,{\rm leading\,automatically\,to}\,\,u^{0}_{1\alpha}=0.$ (18)
@@
$V^{LD}_{relaxed}(\eta)=E_{0}-\frac{1}{2}\Pi_{\tau\alpha}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}\Pi_{\tau^{\prime}\alpha^{\prime}}+\Xi_{\mu\nu}\eta_{\mu\nu}-\Pi_{\tau\alpha}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}Q_{\tau^{\prime}\alpha^{\prime};\mu\nu}\eta_{\mu\nu}+\frac{1}{2}(A_{\mu\nu,\mu^{\prime}\nu^{\prime}}-Q_{\tau\alpha;\mu\nu}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}Q_{\tau^{\prime}\alpha^{\prime};\mu^{\prime}\nu^{\prime}})\eta_{\mu\nu}\eta_{\mu^{\prime}\nu^{\prime}}$ (19)
@@
Note that even if the reference configuration is not exactly the total energy minimum, due to incomplete or unconverged numerical relaxation, eqs. 22 and 18 can in principle be applied (in this order as eq. 18 involves the strain) to further relax the reference primitive cell (these corrections are non-zero only if $\Pi\neq 0$). This can happen in crystals of low symmetry, where space group symmetry constraints do not imply $\Pi=0$. In cases where $\Pi=0$ due to symmetry, such as in Ge, there might additionally be a residual stress because the lattice constant relaxation was not complete. This residual stress is however available at the end of the DFT-SCF relaxation and should be added to the expression for $\sigma_{residual}$ if the equilibrium strain in eq. 22 is needed.

### v2

@@ Abstract
PROGRAM SUMMARY
Program Title: ALATDYN
CPC Library link to program files: (to be added by Technical Editor)
Developer’s repository link:
https://github.com/KeivanS/Anharmonic-lattice-dynamics/tree/main
Licensing provisions: GPLv3
Programming language: FORTRAN90
Nature of problem:
The ALATDYN suite of codes develops a lattice dynamical model of a crystalline solid. The FOCEX code extracts the model parameters from supercell calculations data on forces versus position and calculates the phonon spectrum, elastic constants and thermodynamic properties within the quasi-harmonic approximation. The SCOP8 code goes beyond QHA and implements the self-consistent phonon theory to minimize the free energy with respect to the strain tensor, atomic positions and harmonic force constants, and thus obtains the state of equilibrium at the given temperature along with the effective phonon bands structure. The THERMACOND code uses the cubic force constants and the crystal symmetries to solve the phonon Boltzmann equation (PhBE) efficiently and deduce the thermal conductivity. Finally, ANFOMOD uses the extracted force constants to perform a molecular dynamics simulation in a supercell.
Solution method:
Force constants are obtained from a singular value decomposition (or the ridge regression) method. PhBE is solved by first setting up the collision matrix and effectively inverting it using the conjugate-gradients method.
Additional comments including restrictions and unusual features:
This code has the ability to extract force constants up to 8th order provided enough force-displacement data is provided. SCOP8 can thus use them to potentially predict phase changes. Coupling of phonons to magnetic or orbital degrees of freedom is an extension we plan to add to the FOCEX and SCOP8 codes in the future.
@@
The state of thermal equilibrium (atomic positions and crystal shape and size) as well as the effective harmonic force constants can be predicted at any temperature and volume (or pressure)
@@
Can also perform molecular dynamics simulations and calculate atomic trajectories and heat current using the extracted anharmonic force field. This is the more accurate alternative to solving BE, as it treats anharmonicity and mass disorder beyond perturbation theory.
@@ 1 Introduction
The theory of lattice dynamics, based essentially on the physics of the harmonic oscillator has shaped the science behind the thermophysical properties of materials. Nowadays, with the advances in computer software and hardware, calculations of the electronic ground state properties of materials, based on the density functional theory has become routine. Total energy and forces on atoms are an output of such theories, and this has allowed us [1] to develop methodologies to compute force derivatives, or force constants (FCs), in order to obtain the parameters of an anharmonic lattice dynamics model. Many softwares are nowadays available to compute force constants of crystalline solids. Among them, we can name ALAMODE [2, 3], ShengBTE [4], AlmaBTE [5], Phonopy [6], SSCHA [7] , Hiphive [8], TDEP [9, 10], Phonon [11], Phon [12] … and many others. Here, we are presenting a new set of codes based on our original paper [1] to compute thermophysical properties of anharmonic crystals. These codes are similar in capabilities but different in the way the FCs are extracted and processed. Our code to calculate the thermal conductivity (THERMACOND) uses full symmetry of the crystal to compute the non-equilibrium distribution function more efficiently. Codes to compute the free energy and the state of equilibrium at high temperatures, as well as the computation of mechanical properties are relatively novel[13, 14]. In addition, we provide codes (ANFOMOD) to perform molecular dynamics simulations and calculate the heat current with the extracted polynomial force field. In what follows, we describe the methodology and capabilities of the code to extract force constants, FOCEX, which stands for FOrce Constants EXtraction. The FCs are extracted from a set of displacement-force data on atoms in one or many supercells. They can be used as input to the code ANFOMOD (ANharmonic FOrce MOlecular Dynamics) to perform MD simulations and calculate the heat current and thermal conductivity from the exact Green-Kubo formula, thereby incorporating anharmonicity and defect scattering beyond perturbation theory. Our other codes, which take as input the extracted force constants, compute the thermal conductivity (THERMACOND) and the state of thermal equilibrium, including structural phase changes, at finite temperatures (SCOP8). They will be described in other upcoming papers.
@@ 2.2 Invariance under arbitrary translations and rotations
The potential energy of a solid remains the same under arbitrary translations and rotations of that solid. For any given atomic displacement $u$, the potential energy associated with the translation of the whole system by an arbitrary vector $c$ is $V(u+c)$ and it must satisfy $V(u+c)=V(u)$ and $F(u+c)=F(u)$, where $u$ are the dynamical variables and $c$ is the arbitrary constant shift vector. Identifying terms order by order we arrive at the following equations [15]:
@@
Similarly, the force on atoms and the total potential energy of the crystal should not change under arbitrary rotation of the crystal. This leads to the following constraints on FCs [15]:
@@
We see that rotational invariance relations relate the first to the second-order terms, the second to the third-order terms and the third to the fourth-order terms, respectively. This implies that if the expansion of the potential energy is truncated after the fourth order terms, we must start with the fourth order terms, and the application of the rotational invariance rules will give us constraints on the third and lower-order FCs. For more details on the proof regarding the translation and rotational invariance relations, we refer the reader to the article by Liebfried and Ludwig [15].
@@
In addition to these symmetries, it is important to note that elastic constants are subject to a number of constraints that enforces them to be symmetric under the exchange of Voigt indices. Huang discusses these invariances[16], which provide an additional 15 set of constraint equations for low symmetry crystals with internal degrees of freedom in the primitive cell [17, 18]]. These relations are explicitly written in section 3 where we define the elastic constants.
@@ 2.3 Invariance under point or space group symmetry
Other symmetry operations such as rotations or mirror symmetries, elements of the space group also leave the potential energy of the crystal unaltered [19]. In particular, force constants must be invariant under translations of lattice vectors, implying the following relations:
@@ 2.5 Phonon dispersion and their density of states
$D_{\tau,\tau^{\prime}}^{\alpha\beta}(k)\,e_{\tau^{\prime}\beta,\lambda}(k)=\omega_{\lambda}^{2}(k)\,e_{\tau\alpha,\lambda}(k)$ (9)
$\sum_{\lambda}e_{\tau\alpha,\lambda}(k)^{*}e_{\tau^{\prime}\beta,\lambda}(k)=\delta_{\alpha\beta}\delta_{\tau\tau^{\prime}}\,\,\,\,and\,\,\,\,\\ \sum_{\tau\alpha}e_{\tau\alpha,\lambda}(k)^{*}e_{\tau\alpha,\lambda^{\prime}}(k)=\delta_{\lambda\lambda^{\prime}}$ (10)
@@
The phonon density of states is calculated using both the tetrahedron method in the full first Brillouin zone (FBZ) and the Gaussian-smearing method in the irreducible Brillouin zone (IBZ). The latter will be inaccurate unless a large number of k-points are used in the sampling of the IBZ, while the former is almost converged for a moderate grid size on the order of 20 k-mesh per direction.
@@ 2.6 Non-analytic correction for ionic dielectrics
For such systems, the input to the file dielectric.params which contains the electronic contribution to the dielectric matrix $\epsilon_{\infty}$ and the Born effective charges $Z_{\tau}^{\alpha\beta}$ of every atom $\tau$ in the primitive cell, must be properly initialized. These numbers are obtained from a separate DFPT calculation[20, 21]. The contribution of the ionic charges on the force constants is long-ranged and cannot be captured by the small supercell. It is usually calculated using the Ewald sum method [22, 23, 20], which is formed from the the sum of short-range real space component and a long-range reciprocal-space component. In FOCEX, once the harmonic force constants are fitted, we calculate the dynamical matrix on a grid of reciprocal lattice vectors of the supercell inside the Wigner-Seitz cell of the primitive lattice, including its boundaries, albeit with a weight inversely proportional to the number of images that also fall on the Brillouin zone boundaries. This corresponds to the Fourier transform of the real-space force constants. Phonon frequencies on this grid, $D(G_{\rm sc})$, are exact. From these dynamical matrices, we subtract the non-analytical contribution[20], and Fourier transform back to real space, in-order to presumably obtain a short-ranged set of force constants, with a range confined to the Wigner-Seitz cell of the supercell. Then at an arbitrary wavevector $q$, we use Fourier interpolation to calculate the dynamical matrix from the latter set of force constants, to which the non-analytical part that was subtracted earlier, is added back (see also [24]).
@@ 3 Some notes on Elastic constants
It is possible to extract the elastic constants from the knowledge of the force constants[25]. Below, we outline a simpler method and provide the final formulas.
@@
While this is valid strictly at zero temperature, this discussion can be extended to finite temperatures using the self-consistent phonon theory[26] where the harmonic force constant $\Phi$ is replaced by an effective temperature-dependent one[27, 28]:
$\Phi\rightarrow K(T)=\Phi+\Psi S+\chi(SS+\langle yy\rangle)/2$ (14)
. In what follows, we will derive the (harmonic) elastic properties, strictly speaking, at zero temperature, but the above formula may be used as an extension to finite temperatures, by replacing every occurrence of $\Phi$ in the following by $K(T)$.
@@
Compared to the method discussed by Wallace [25], here, we are proposing an equivalent but simpler method to obtain elastic constants. We first write the harmonic potential energy in eq. 13, phenomenologically, up to second powers of internal coordinates $u^{0}$ and strain $\eta$ as:
$V^{LD}(\eta,u^{0})=E_{0}+\Pi_{\tau\alpha}u^{0}_{\tau\alpha}+\Xi_{\mu\nu}\eta_{\mu\nu}+\\ \frac{1}{2}\left[\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}\,u^{0}_{\tau\alpha}u^{0}_{\tau^{\prime}\alpha^{\prime}}+A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}\,\eta_{\alpha\beta}\eta_{\alpha^{\prime}\beta^{\prime}}+2Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}\,u^{0}_{\tau\alpha}\eta_{\alpha^{\prime}\beta^{\prime}}\right]$ (15)
@@
$\Xi_{\alpha\beta}$ $=$ $\frac{1}{N}\sum_{R\tau}\Pi_{\tau\alpha}(R+\tau)^{\beta}=\sum_{\tau}\Pi_{\tau\alpha}\tau^{\beta}$
$\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}$ $=$ $\frac{1}{N}\sum_{R,R^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}\,\,\,\,{\rm with}\,\,\,\,\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\phi_{\tau^{\prime},\tau}^{\alpha^{\prime}\alpha}$
$Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}$ $=$ $\frac{1}{N}\sum_{RR^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R^{\prime}+\tau^{\prime})^{\beta^{\prime}}=\frac{1}{N}\sum_{RR^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R^{\prime}+\tau^{\prime}-R-\tau)^{\beta^{\prime}}$
$A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}$ $=$ $\frac{1}{N}\sum_{R\tau,R^{\prime}\tau^{\prime}}\Phi_{R\tau,R^{\prime}\tau^{\prime}}^{\alpha,\alpha^{\prime}}(R+\tau)^{\beta}(R^{\prime}+\tau^{\prime})^{\beta^{\prime}}\,\,{\rm with}\,\,A_{\alpha\beta,\alpha^{\prime}\beta^{\prime}}=A_{\alpha^{\prime}\beta^{\prime},\alpha\beta}$
${\rm ASR}$ ==>= > $\sum_{\tau}\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\sum_{\tau^{\prime}}\phi_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}=0$
${\rm ASR}$ ==>= > $\sum_{\tau}Q_{\tau\alpha;\alpha^{\prime}\beta^{\prime}}=0$
@@
$u^{0}_{\tau\alpha}(\eta)=-\Gamma_{\tau,\tau^{\prime}}^{\alpha\alpha^{\prime}}(\Pi_{\tau^{\prime}\alpha^{\prime}}+Q_{\tau^{\prime}\alpha^{\prime};\mu\nu}\,\eta_{\mu\nu})\\ {\rm with}\,\,\,\Gamma_{1,\tau^{\prime}}^{\alpha\alpha^{\prime}}=\Gamma_{\tau,1}^{\alpha\alpha^{\prime}}=0\,\,{\rm leading\,automatically\,to}\,\,u^{0}_{1\alpha}=0.$ (16)
@@
$V^{LD}_{relaxed}(\eta)=E_{0}-\frac{1}{2}\Pi_{\tau\alpha}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}\Pi_{\tau^{\prime}\alpha^{\prime}}+\Xi_{\mu\nu}\eta_{\mu\nu}-\Pi_{\tau\alpha}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}Q_{\tau^{\prime}\alpha^{\prime};\mu\nu}\eta_{\mu\nu}\\ +\frac{1}{2}(A_{\mu\nu,\mu^{\prime}\nu^{\prime}}-Q_{\tau\alpha;\mu\nu}\Gamma_{\tau\alpha,\tau^{\prime}\alpha^{\prime}}Q_{\tau^{\prime}\alpha^{\prime};\mu^{\prime}\nu^{\prime}})\eta_{\mu\nu}\eta_{\mu^{\prime}\nu^{\prime}}$ (17)
@@
Note that even if the reference configuration is not exactly the total energy minimum, due to incomplete or unconverged numerical relaxation, eqs. 20 and 16 can in principle be applied (in this order as eq. 16 involves the strain) to further relax the reference primitive cell (these corrections are non-zero only if $\Pi\neq 0$). This can happen in crystals of low symmetry, where space group symmetry constraints do not imply $\Pi=0$. In cases where $\Pi=0$ due to symmetry, such as in Ge, there might additionally be a residual stress because the lattice constant relaxation was not complete. This residual stress is however available at the end of the DFT-SCF relaxation and should be added to the expression for $\sigma_{residual}$ if the equilibrium strain in eq. 20 is needed.